OpenStudy (anonymous):

in the early months of some year, one site added 02 million new accounts every day.at this rate, how many days would be needed to add new 12 million accounts? it takes__days to add 12million accounts

OpenStudy (anonymous):

is 02 milllion equal to a) 2 million or b) 0.2 million? if a) 2 000 000 * x = 12 000 000 now you can divide both side with 1 000 000: 2*x = 12 if b) 0.2 million is 200 000 200 000 * x = 12 000 000 now you can divide both sides with 100 000 2*x=120 (in both cases x is the number of days)